OpenStudy (anonymous):

Maria received a gift of cash from her grandmother. She put two-thirds of the money in her savings account, spent twenty-five dollars on a pair of jeans, and spent half the remaining money on a DVD. If Maria had eighteen dollars left over, how much money did she receive from her grandmother?

OpenStudy (anonymous):

after spending half on DVD she msut have been left with another half part which is 18 here.. so after spending 25dollars on jeans she was left with 18/(1/2) =18*2 =36

OpenStudy (anonymous):

before spending for jeans she must have 36+25 =61 which is (1-2/3) of what her grandmother gifted.. and so on..

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

welcome..

OpenStudy (anonymous):

its that all

OpenStudy (anonymous):

no more steps

OpenStudy (anonymous):

@matricked

OpenStudy (anonymous):

no you have to find 1-2/3 and then divide 61 by (1-2/3)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is it 183 @matricked

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

ok thaank u

OpenStudy (anonymous):

yw